DVITYPE(1) DVITYPE(1) NAME dvitype−translateadvifileforhumans SYNOPSIS dvitypedviname[.dvi] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. ThedvitypeprogramtranslatesaDVI(DeViceIndependent)fileoutputby(forexample)tex(1) orgftodvi(1),toafilethathumanscanread.ItalsoservesasaDVIfile-validatingprogram(i.e., ifdvitypecanreadit,it’scorrect)andasanexampleofaDVI-readingprogramforfuturedevice drivers. Theoutputfilecanincludeallcommands,justtheimportantones,ornoneatall(inwhichcase onlyerrorsarereported).Asubintervalofpagesmaybeselectedfortransliteration;themagnificationandresolutionofthe‘‘outputdevice’’maybechanged;andsoon.Alloptionsarespecified withanon-linedialog. The.dviextensionissuppliedifomittedfromdviname.Theoutputgoestostdout. ENVIRONMENT TheenvironmentvariableTEXFONTSisusedtosearchfortheTFMfilesusedintheDVIfile. Seetex(1)forthedetailsofthesearching.IfTEXFONTSisnotset,itusesthesystemdefault. SEEALSO gftype(1),pktype(1). DonaldE.Knuth,TEXware. BUGS Theinteractivedialogshouldbereplacedbycommand-lineoptions,aswithgftype. AUTHORS DonaldE.Knuthwrotetheprogram.ItwaspublishedaspartoftheTEXwaretechnicalreport, availablefromtheTEXUsersGroup.HowardTrickeyandPavelCurtisoriginallyporteditto Unix. Web2C7.32x 19December1994 1